# Physics-Driven Neural Compensation For Electrical Impedance Tomography

### v1

@@ Abstract
Index Terms:
implicit neural representation, electrical impedance tomography, unsupervised learning.
I Introduction
@@
The rest of this paper is organized as follows. Section II briefly revisits the mathematical model of EIT and reviews existing neural field works for image reconstruction. Section III proposes a EIT reconstruction method with neural compensation strategy. Experimental results are systematically presented in Section IV, followed by a comprehensive analysis of outcomes. Finally, Sections V and VI offer discussions and conclusions, respectively.
II Background
II-A Mathematical Model of EIT
@@ II-B Neural Field
III Physics Driven Neural Compensation
In this section, we will first provide a brief introduction to the PhyNC framework. Then, the details will be introduced, including sensitivity based level mapping in Section III-B, hybrid representations in Section III-C, and frequency regularization in Section III-D.
III-A Overview
@@
III-B Sensitivity-Based Level Mapping
@@
Building on this definition, we compared PhyNC’s initial sensitivity distribution before iteration with that of Hash [60] to illustrate the impact of neural compensation. In Fig.2, we visualize their differences within an uniform conductivity domain. Hash, which encodes all coordinates uniformly, demonstrates markedly lower sensitivity in the central region due to limitations in its underlying physics model. In contrast, PhyNC achieves a more balanced sensitivity distribution through its level mapping process, boosting sensitivity precisely where the physics model traditionally underperforms. However, it is worth noting that as the neural representation of the conductivity $\sigma(\mathbf{x})$ evolves during iteration, $\partial\sigma/\partial\mathbf{x}$ (and thus overall sensitivity) may adaptively redistribute. While PhyNC’s initialization mitigates EIT’s inherent sensitivity imbalance, the fidelity of its final reconstructions (especially in low-sensitivity regions) must be empirically validated, as discussed in Section IV.
III-C Hybrid Representation
@@
III-D Frequency Regularization for Neural Field Optimization
@@
TABLE I: Quantitative Comparison of Our Method versus baselines with Simulated Data Featuring Various Geometric Shape Combinations
@@
IV Experiments and Results
IV-A Experimental Setup
IV-A1 Data
@@ IV-A3 Implementation Details
IV-B Detail recovery and Anti-artifacts by PhyNC
We first evaluated the performance of PhyNC using simulated data. Specifically, we tested the method on reconstructions of three basic geometric shapes: a circle, a rectangle, and a triangle. The reconstruction results are shown in Fig. 3(a). PhyNC effectively reconstructs each object’s shape, size, and position, with notable accuracy in capturing the right angles of rectangles and the acute angles of triangles. Moreover, the proposed method shows superior image quality compared to reference methods, as confirmed by quantitative evaluation metrics with results summarized in Table I.
@@ IV-B Detail recovery and Anti-artifacts by PhyNC
IV-C Sensitivity compensation and Reconstruction robustness by PhyNC
@@
IV-D PhyNC for experimental data
@@
IV-E Ablation study and Representation analysis
@@
IV-F Trade-Off Between FEM and Neural Field
@@
These observations underscore the importance of selecting an optimal FEM mesh resolution. Notably, with a moderately fine mesh ($N_{g}=3,154$), most conventional methods begin to decline in performance. In contrast, our proposed PhyNC method continues to improve, reaching peak performance at this discretization level. To quantitatively assess each method’s resilience to increasing ill-posedness, we performed two-stage linear fittings on standard image quality metrics—PSNR, SSIM, and LPIPS—to identify a ”limiting number of nodes” for each approach. (It is important to note that these limiting values are derived solely from the fitting process and were not directly tested with meshes corresponding to those exact node counts.)

### v2

@@ Abstract
1 Introduction
@@
The rest of this paper is organized as follows. Section 2 briefly revisits the mathematical model of EIT and reviews existing neural field works for image reconstruction. Section 3 proposes a EIT reconstruction method with neural compensation strategy. Experimental results are systematically presented in Section 4, followed by a comprehensive analysis of outcomes. Finally, Sections 5 and 6 offer discussions and conclusions, respectively.
2 Background
2.1 Mathematical Model of EIT
@@ 2.2 Neural Field
3 Physics Driven Neural Compensation
In this section, we will first provide a brief introduction to the PhyNC framework. Then, the details will be introduced, including sensitivity based level mapping in Section 3.2, hybrid representations in Section 3.3, and frequency regularization in Section 3.4.
3.1 Overview
@@
3.2 Sensitivity-Based Level Mapping
@@
Building on this definition, we compared PhyNC’s initial sensitivity distribution before iteration with that of Hash [60] to illustrate the impact of neural compensation. In Fig.2, we visualize their differences within an uniform conductivity domain. Hash, which encodes all coordinates uniformly, demonstrates markedly lower sensitivity in the central region due to limitations in its underlying physics model. In contrast, PhyNC achieves a more balanced sensitivity distribution through its level mapping process, boosting sensitivity precisely where the physics model traditionally underperforms. However, it is worth noting that as the neural representation of the conductivity $\sigma(\mathbf{x})$ evolves during iteration, $\partial\sigma/\partial\mathbf{x}$ (and thus overall sensitivity) may adaptively redistribute. While PhyNC’s initialization mitigates EIT’s inherent sensitivity imbalance, the fidelity of its final reconstructions (especially in low-sensitivity regions) must be empirically validated, as discussed in Section 4.
3.3 Hybrid Representation
@@
3.4 Frequency Regularization for Neural Field Optimization
@@
Table 1: Quantitative Comparison of Our Method versus baselines with Simulated Data Featuring Various Geometric Shape Combinations
@@
4 Experiments and Results
4.1 Experimental Setup
4.1.1 Data
@@ 4.1.3 Implementation Details
4.2 Detail recovery and Anti-artifacts by PhyNC
We first evaluated the performance of PhyNC using simulated data. Specifically, we tested the method on reconstructions of three basic geometric shapes: a circle, a rectangle, and a triangle. The reconstruction results are shown in Fig. 3(a). PhyNC effectively reconstructs each object’s shape, size, and position, with notable accuracy in capturing the right angles of rectangles and the acute angles of triangles. Moreover, the proposed method shows superior image quality compared to reference methods, as confirmed by quantitative evaluation metrics with results summarized in Table 1.
@@ 4.2 Detail recovery and Anti-artifacts by PhyNC
4.3 Sensitivity compensation and Reconstruction robustness by PhyNC
@@
4.4 PhyNC for experimental data
@@
4.5 Ablation study and Representation analysis
@@
4.6 Trade-Off Between FEM and Neural Field
@@
These observations underscore the importance of selecting an optimal FEM mesh resolution. Notably, with a moderately fine mesh ($N_{g}=3,154$), most conventional methods begin to decline in performance. In contrast, our proposed PhyNC method continues to improve, reaching peak performance at this discretization level. To quantitatively assess each method’s resilience to increasing ill-posedness, we performed two-stage linear fittings on standard image quality metrics—PSNR, SSIM, and LPIPS—to identify a "limiting number of nodes" for each approach. (It is important to note that these limiting values are derived solely from the fitting process and were not directly tested with meshes corresponding to those exact node counts.)